Yumiko has just had her first child. How much money should she invest at 4% APR compouned monthly to have $50,000 18 years from now?
Answer by Alan3354(69443)

Yumiko has just had her first child. How much money should she invest at 4% APR compouned monthly to have $50,000 18 years from now?
----------------
A = amount to invest
FV = future value in 18 yrs
18 yrs = 216 months
---------
FV = A*(1 + 0.04/12)^216
50000 = A*(1 + 0.01/3)^216
A = 50000/((1+0.01/3)^216
A =~ $24366.77